OpenStudy (anonymous):

3 questions 1.Create a set A and a subset of A. Name the subset S. Find S'. 2.In your own words, explain what the vertical line test is and how it is used 3.A = {2, 4, 6} and let B = {circle, square}. Find A X B

OpenStudy (anonymous):

2. In \(my\) own words: For any point \(x\), there may be no more than 1 value \(y\) that satisfies \(y=f(x)\). The vertical line test is used to test whether or not any relation between \(x\) and \(y\) is also a function of \(x\).

OpenStudy (anonymous):

3. \(\{(2,~circle),~(2,~square),~(4,~circle),~(4,~square),~(6,~circle),~(6,~square)\}\)

OpenStudy (anonymous):

1. \(S' = A - S\) or \(S' = A \setminus S\) They are identical.